# Modular analogs of character formulas and multiplicities of induced representations

### v1

@@ A.1. Introduction
    and also review a criterion due to Vignéras. Our necessary criterion is very far from these sufficient criteria and the problem of precisely determining when $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ is self-projective or self-injective seems difficult.
When the sufficient criteria for these two points are both met, then one can conclude that the socle and cosocle of $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ are multiplicity free. One simple case where the criteria are both met is when $H$ has prime-to-$p$ order and $\rho$ is a character. This case recovers [Zha24, Corollaries 5.2 and 5.4] about multiplicity-freeness of modular Gelfand–Graev modules (indeed, as noted in [Zha24, Remark 5.3], these statements are vacuous in the equal-characteristic case $\ell\mid q$, while the relevant group $H$ has prime-to-$p$ order in the unequal characteristic case $\ell\nmid q$). This also recovers the applications of [Zha24] used in [BGGG+23, Section 4.1].
A.1.1. Further comments
To avoid possible confusion about how errors in [Zha24] affect the results here, we have made this appendix completely independent of [Zha24] and do not assume that the reader is familiar with [Zha24]. We have attempted to reproduce corrected versions of the most important results of [Zha24] here, so that the reader interested in those results need only look at this paper (although we note that [Zha24, Sections 2, 3, and 6] are not affected by the errors that we identify). For the most part, we only deal with finite groups and have not attempted to extend the results to infinite compact groups, which are the subject of [Zha24, Corollary 1.10 and Sections 4.3, 5.2]. However, in Section A.4, we reproduce a proof for the $\ell=p$ case of [Zha24, Corollary 1.10] that we learned from Dipendra Prasad. We do not know whether the $\ell\neq p$ case of [Zha24, Corollary 1.10] is true because the methods we develop here are not sufficient to determine whether self-injectivity holds.
A.2. Endomorphisms of induced representations
@@
A.2.1. Reduction modulo $p$ on endomorphisms of inductions
@@
${\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}\rho)}}\to\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$
@@
The first statement is simply a combination of Lemma A.2 and Mackey’s formula (A.1). For $n=1$, (A.2) simplifies to
@@
Corollary A.4.
Let $\rho:H\to\operatorname{GL}_{n}(\mathcal{O})$ be an $\mathcal{O}$-lattice in a representation $\rho_{K}$ over $K$, and suppose that $\mathrm{Ind}_{H}^{G}\rho_{K}$ is multiplicity free and that (A.2) holds. Then $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$ is commutative.
@@
Since $\mathrm{Ind}_{H}^{G}(\rho_{K})$ is multiplicity free, $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}\rho_{K})$ is commutative. Hence the subquotient ring $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}\rho)}$} is a also commutative. By Proposition A.3, $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$ is isomorphic to $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}\rho)}$}, so it is also commutative. ∎
A.2.2. Examples
We give some examples where (A.2) fails and where the mod-$p$ twisted Hecke algebra is noncommutative even where the characteristic zero version is commutative.
Example A.5.
Let $p$ be an odd prime, $\mathcal{O}=\mathbb{Z}_{p}[\zeta_{p}]$, $G=\operatorname{GL}_{2}(\mathbb{F}_{p})$, $H$ be the subgroup of upper-triangular unipotent matrices, and $\rho:H\to\mathcal{O}^{\times}$ be a nontrivial character. By (A.1), the dimension of $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}\rho)$ is equal to
@@ A.2.2. Examples
We now come to Example A.1 from the introduction. We first give a simple criterion for $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}\rho)$ to be commutative.
Lemma A.6.
Let $N$ be a normal subgroup of $G$ and let $\rho:N\to K^{\times}$ be a character. Since $N$ is normal, $\rho^{s}$ is also a character of $N$ for all $s\in G$. Let $H=\{s\in G\ |\ \rho^{s}=\rho\}$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)$ is commutative (and $\mathrm{Ind}_{N}^{G}\rho$ is multiplicity free).
@@
Since $N$ is normal, $\G/N$$N\backslash G$/N=G/N. By (A.1),
@@
We also have a simple criterion for the socle and cosocle of $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ to have multiplicity.
Lemma A.7.
Let $G$ be a group with a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Then for every character $\rho:N\to K^{\times}$, the induction $\mathrm{Ind}_{N}^{G}{\overline{\rho}}$ is semisimple and is not multiplicity free.
@@
Since $N$ is a $p$-group, ${\overline{\rho}}=1$ is the trivial character. Since $N$ is normal, $\mathrm{Ind}_{N}^{G}1$ is the inflation of the regular representation $k[G/N]$. Because $p\nmid\#(G/N)$, the usual formula for the regular representation in characteristic zero,
@@
Corollary A.8.
Let $G$ be a group that has a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Let $\rho:N\to K^{\times}$ be a character and let $H$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{Ind}_{N}^{G}\rho$ is multiplicity free but $\mathrm{Ind}_{N}^{G}{\overline{\rho}}$ is not.
A minimal example of a group satisfying the hypothesis of Corollary A.8 is $G=D_{30}$, the dihedral group of order $30$, with $p=5$ as in Example A.1. The $5$-Sylow subgroup $N$ is normal and for every nontrivial $\rho:N\to K^{\times}$, the stabilizer $H$ of $\rho$ is the centralizer $C_{G}(N)$ of $N$, which is cyclic of order $15$. Hence $[H:N]=3<4$, and Corollary A.8 applies.
A.3. Self-projectivity of induced representations
@@
A.3.1. Projective covers and injective envelopes
@@
A.3.2. Self-projectivity and self-injectivity
@@
There is a dual lemma about self-injectivity, whose proof is almost the exact dual of the proof of Lemma A.10.
Lemma A.11.
@@
A.3.3. Failure of self-projectivity and self-injectivity of certain inductions
@@
Lemma A.12.
Let $G$ be a finite group and let $H$ be a non-normal subgroup of $G$. Let $M=\mathrm{Ind}_{H}^{G}1$ be the induction of the trivial representation of $H$. Suppose that $k[G]$ is a projective cover (resp. injective envelope) of $M$. Then $M$ is not self-projective (resp. self-injective).
@@
Suppose that $k[G]$ is a projective cover of $M$. By Lemma A.10, to show that $M$ is not self-projective, it is enough to show that
@@
Now suppose that $k[G]$ is an injective envelope of $M$. By Lemma A.11, to show that $M$ is not self-injective, it is enough to show that
@@
Proposition A.13.
If $G$ is a $p$-group and $H$ is a non-normal subgroup of $G$, then $\mathrm{Ind}_{H}^{G}1$ is neither self-projective nor self-injective.
@@
By the previous lemma, it is enough to show that $k[G]$ is a projective cover and injective envelope of $\mathrm{Ind}_{H}^{G}1$. Inducing the surjection $k[H]\twoheadrightarrow k$ shows that $k[G]$ surjects onto $\mathrm{Ind}_{H}^{G}1$, so the projective cover of $\mathrm{Ind}_{H}^{G}1$ is a direct summand of $k[G]$. Similarly, inducing the injective $k\hookrightarrow k[H]$ shows that $\mathrm{Ind}_{H}^{G}1$ can be embedded in $k[G]$, so the injective envelope of $\mathrm{Ind}_{H}^{G}1$ is a direct summand of $k[G]$. But, since $k[G]$ is local, it has no proper direct summands. ∎
A.3.4. Criteria for self-projectivity of inductions
@@
Lemma A.14.
Let $G$ be a finite group with a subgroup $H$ and let $M$ be a $k[H]$-module. Then $\mathrm{Ind}_{H}^{G}M$ is self-projective if any one of the following holds:
@@
If (1) holds, then $\mathrm{Ind}_{H}^{G}M$ is self-projective by a Frobenius-reciprocity argument (see [Zha24, Corollary 4.2(ii)]). If (2) holds, then $k[H]$ is a semisimple ring, so $M$ is a projective $k[H]$-module and (1) holds. If (3) holds, then $\mathrm{Ind}_{H}^{G}M|_{H}$ is a semisimple $k[H]$-module by Mackey’s formula, so every $k[H]$-module NN is $\mathrm{Ind}_{H}^{G}M|_{H}$-projective and (1) holds again. ∎
The following criterion is a slight generalization of [Vig01, Lemma 3.1], in that we show that the assumption there that $M_{\rho}$ is a direct factor of $M|_{H}$ is unnecessary.
Lemma A.15.
Let $G$ be a finite group with a subgroup $H$. Let $\rho$ be a simple $k[H]$-module and let $M=\mathrm{Ind}_{H}^{G}\rho$. Let $M_{\rho}$ be the largest $k[H]$-submodule of $M|_{H}$ that is isomorphic to a direct sum of copies of $\rho$, and let $M^{\rho}=(M|_{H})/M_{\rho}$. If no subquotient of $M^{\rho}$ is isomorphic to $\rho$, then $M$ is self-projective.
@@
$\mathrm{End}_{k[G]}(M)=\mathrm{Hom}_{k[H]}(\rho,M|_{H})=\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$
@@
$\mathrm{Hom}_{k[G]}(P,M)=\mathrm{Hom}_{k[H]}(P_{\rho},M|_{H}).$
Since $\mathrm{Hom}_{k[H]}(P_{\rho},M^{\rho})=0$ by Lemma A.9, it follows that
$\mathrm{Hom}_{k[H]}(P_{\rho},M|_{H})=\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho}).$
But every homomorphism $P_{\rho}\to M_{\rho}$ factors through the cosocle $\rho$ of $P_{\rho}$, so $\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho})\cong\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$. Altogether, this shows that $\mathrm{End}_{k[G]}(M)$ and $\mathrm{Hom}_{k[G]}(P,M)$ are isomorphic, so $M$ is self-projective by Lemma A.10
Vignéras uses this criterion to show [Vig01, Corollary 5.2] that $\mathrm{Ind}_{P}^{G}\sigma$ is self-projective if $G=\mathbf{G}(F)$ is the $F$-rational points of a reductive group for a field $F$ with $\mathrm{char}(F)\neq\mathrm{char}(k)$, $P=MU$ is a parabolic subgroup with Levi $M$ and unipotent radical $U$, and $\sigma$ is a cuspidal representation of $M$.
A.4. Quaternion division algebras

### v2

# Corrigendum to “Modular Gelfand pairs and multiplicity-free representations”

@@ 1. Introduction
    and also review a criterion due to Vignéras. Our necessary criterion is very far from these sufficient criteria and the problem of precisely determining when $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ is self-projective or self-injective seems difficult.
When the sufficient criteria for these two points are both met, then one can conclude that the socle and cosocle of $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ are multiplicity free. One simple case where the criteria are both met is when $H$ has prime-to-$p$ order and $\rho$ is a character. This case recovers [Zha24, Corollaries 5.2 and 5.4] about multiplicity-freeness of modular Gelfand–Graev modules (indeed, as noted in [Zha24, Remark 5.3], these statements are vacuous in the equal-characteristic case $\ell\mid q$, while the relevant group $H$ has prime-to-$p$ order in the unequal characteristic case $\ell\nmid q$). This also recovers the applications of [Zha24] used in [BGGG+23, Section 4.1].
To avoid possible confusion about how errors in [Zha24] affect the results here, we have made this article completely independent and do not assume that the reader is familiar with [Zha24]. We have attempted to reproduce corrected versions of the most important results of [Zha24] here, so that the reader interested in those results need only look at this paper (although we note that [Zha24, Sections 2, 3, and 6] are not affected by the errors that we identify). For the most part, we only deal with finite groups and have not attempted to extend the results to infinite compact groups, which are the subject of [Zha24, Corollary 1.10 and Sections 4.3, 5.2]. However, in Section 4, we reproduce a proof for the $\ell=p$ case of [Zha24, Corollary 1.10] that we learned from Dipendra Prasad. We do not know whether the $\ell\neq p$ case of [Zha24, Corollary 1.10] is true because the methods we develop here are not sufficient to determine whether self-injectivity holds.
Acknowledgements
R. Z. is extremely grateful to his coauthors for identifying and communicating the errors and for agreeing to work together on this note. He also thanks Karol Koziol for identifying some of the errors and pointing out Example 2.4; Ofer Gabber for explaining the equivalence of self-projectivity and almost-projectivity when there are enough projectives; and Dipendra Prasad for pointing out how to recover Corollary 4.1.
2. Endomorphisms of induced representations
@@
2.1. Reduction modulo $p$ on endomorphisms of inductions
@@
${\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}(\rho))}}\to\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}({\overline{\rho}}))$
@@
The first statement is simply a combination of Lemma 2.1 and Mackey’s formula (2.1). For $n=1$, (2.2) simplifies to
@@
Corollary 2.3.
Let $\rho:H\to\operatorname{GL}_{n}(\mathcal{O})$ be an $\mathcal{O}$-lattice in a representation $\rho_{K}$ over $K$, and suppose that $\mathrm{Ind}_{H}^{G}(\rho_{K})$ is multiplicity-free and that (2.2) holds. Then $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}(\overline{\rho}))$ is commutative.
@@
Since $\mathrm{Ind}_{H}^{G}(\rho_{K})$ is multiplicity-free, $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}(\rho_{K}))$ is commutative. Hence the subquotient ring $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}(\rho))}$} is a also commutative. By Proposition 2.2, $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}(\overline{\rho}))$ is isomorphic to $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}(\rho))}$}, so it is also commutative. ∎
2.2. Examples
We give some examples where (2.2) fails and where the mod-$p$ twisted Hecke algebra is noncommutative even where the characteristic zero version is commutative.
Example 2.4.
Let $p$ be an odd prime, $\mathcal{O}=\mathbb{Z}_{p}[\zeta_{p}]$, $G=\operatorname{GL}_{2}(\mathbb{F}_{p})$, $H$ be the subgroup of upper-triangular unipotent matrices, and $\rho:H\to\mathcal{O}^{\times}$ be a nontrivial character. By (2.1), the dimension of $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}(\rho))$ is equal to
@@ 2.2. Examples
We now come to Example 1.1 from the introduction. We first give a simple criterion for $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}(\rho))$ to be commutative.
Lemma 2.5.
Let $N$ be a normal subgroup of $G$ and let $\rho:N\to K^{\times}$ be a character. Since $N$ is normal, $\rho^{s}$ is also a character of $N$ for all $s\in G$. Let $H=\{s\in G\ |\ \rho^{s}=\rho\}$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)$ is commutative (and $\mathrm{Ind}_{N}^{G}(\rho)$ is multiplicity-free).
@@
Since $N$ is normal, $\G/N$$N\backslash G$/N=G/N. By (2.1),
@@
We also have a simple criterion for the socle and cosocle of $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ to have multiplicity.
Lemma 2.6.
Let $G$ be a group with a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Then for every character $\rho:N\to K^{\times}$, the induction $\mathrm{Ind}_{N}^{G}(\overline{\rho})$ is semisimple and is not multiplicity free.
@@
Since $N$ is a $p$-group, $\overline{\rho}=1$ is the trivial character. Since $N$ is normal, $\mathrm{Ind}_{N}^{G}(\mathbbm{1})$ is the inflation of the regular representation $k[G/N]$. Because $p\nmid\#(G/N)$, the usual formula for the regular representation in characteristic zero,
@@
Corollary 2.7.
Let $G$ be a group that has a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Let $\rho:N\to K^{\times}$ be a character and let $H$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{Ind}_{N}^{G}(\rho)$ is multiplicity-free but $\mathrm{Ind}_{N}^{G}(\overline{\rho})$ is not.
A minimal example of a group satisfying the hypothesis of Corollary 2.7 is $G=D_{30}$, the dihedral group of order $30$, with $p=5$ as in Example 1.1. The 55-Sylow subgroup NN is normal and for every nontrivial $\rho:N\to K^{\times}$, the stabilizer $H$ of $\rho$ is the centralizer $C_{G}(N)$ of $N$, which is cyclic of order $15$. Hence $[H:N]=3<4$, and Corollary 2.7 applies.
3. Self-projectivity of induced representations
@@
3.1. Projective covers and injective envelopes
@@
3.2. Self-projectivity and self-injectivity
@@
There is a dual lemma about self-injectivity, whose proof is almost the exact dual of the proof of Lemma 3.2.
Lemma 3.3.
@@
3.3. Failure of self-projectivity and self-injectivity of certain inductions
@@
Lemma 3.4.
Let $G$ be a finite group and let $H$ be a non-normal subgroup of $G$. Let $M=\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ be the induction of the trivial representation of $H$. Suppose that $k[G]$ is a projective cover (resp. injective envelope) of $M$. Then $M$ is not self-projective (resp. self-injective).
@@
Suppose that $k[G]$ is a projective cover of $M$. By Lemma 3.2, to show that $M$ is not self-projective, it is enough to show that
@@
Now suppose that $k[G]$ is an injective envelope of $M$. By Lemma 3.3, to show that $M$ is not self-injective, it is enough to show that
@@
Proposition 3.5.
If $G$ is a $p$-group and $H$ is a non-normal subgroup of $G$, then $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ is neither self-projective nor self-injective.
@@
By the previous lemma, it is enough to show that $k[G]$ is a projective cover and injective envelope of $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$. Inducing the surjection $k[H]\twoheadrightarrow k$ shows that $k[G]$ surjects onto $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$, so the projective cover of $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ is a direct summand of $k[G]$. Similarly, inducing the injective $k\hookrightarrow k[H]$ shows that $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ can be embedded in $k[G]$, so the injective envelope of $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ is a direct summand of $k[G]$. But, since $k[G]$ is local, it has no proper direct summands. ∎
3.4. Criteria for self-projectivity of inductions
@@
Lemma 3.6.
Let $G$ be a finite group with a subgroup $H$ and let $M$ be a $k[H]$-module. Then $\mathrm{Ind}_{H}^{G}(M)$ is self-projective if any one of the following holds:
@@
If (1) holds, then $\mathrm{Ind}_{H}^{G}(M)$ is self-projective by a Frobenius-reciprocity argument (see [Zha24, Corollary 4.2(ii)]). If (2) holds, then $k[H]$ is a semisimple ring, so $M$ is a projective $k[H]$-module and (1) holds. If (3) holds, then $\mathrm{Res}_{H}^{G}(\mathrm{Ind}_{H}^{G}(M))$ is a semisimple $k[H]$-module by Mackey’s formula, so every $k[H]$-module NN is $\mathrm{Res}_{H}^{G}(\mathrm{Ind}_{H}^{G}(M))$-projective and (1) holds again. ∎
The following criterion is a slight generalization of [Vig01, Lemma 3.1], in that we show that the assumption there that $M_{\rho}$ is a direct factor of $\mathrm{Res}_{H}^{G}(M)$ is unnecessary.
Lemma 3.7.
Let $G$ be a finite group with a subgroup $H$. Let $\rho$ be a simple $k[H]$-module and let $M=\mathrm{Ind}_{H}^{G}(\rho)$. Let $M_{\rho}$ be the largest $k[H]$-submodule of $\mathrm{Res}_{H}^{G}(M)$ that is isomorphic to a direct sum of copies of $\rho$, and let $M^{\rho}=\mathrm{Res}_{H}^{G}(M)/M_{\rho}$. If no subquotient of $M^{\rho}$ is isomorphic to $\rho$, then $M$ is self-projective.
@@
$\mathrm{End}_{k[G]}(M)=\mathrm{Hom}_{k[H]}(\rho,\mathrm{Res}_{H}^{G}(M))=\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$
@@
$\mathrm{Hom}_{k[G]}(P,M)=\mathrm{Hom}_{k[H]}(P_{\rho},\mathrm{Res}_{H}^{G}(M)).$
Since $\mathrm{Hom}_{k[H]}(P_{\rho},M^{\rho})=0$ by Lemma 3.1, it follows that
$\mathrm{Hom}_{k[H]}(P_{\rho},\mathrm{Res}_{H}^{G}(M))=\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho}).$
But every homomorphism $P_{\rho}\to M_{\rho}$ factors through the cosocle $\rho$ of $P_{\rho}$, so $\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho})\cong\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$. Altogether, this shows that $\mathrm{End}_{k[G]}(M)$ and $\mathrm{Hom}_{k[G]}(P,M)$ are isomorphic, so $M$ is self-projective by Lemma 3.2
Vignéras uses this criterion to show [Vig01, Corollary 5.2] that $\mathrm{Ind}_{P}^{G}(\sigma)$ is self-projective if $G=\mathbf{G}(F)$ is the $F$-rational points of a reductive group for a field $F$ with $\mathrm{char}(F)\neq\mathrm{char}(k)$, $P=MU$ is a parabolic subgroup with Levi $M$ and unipotent radical $U$, and $\sigma$ is a cuspidal representation of $M$.
4. Quaternion division algebras